OpenStudy (anonymous):

please help 1) Solve: 1. 6y + 2 – 4y = -18 Solve each equation for the given variable: 1. 3y + 5x = 4; y

OpenStudy (anonymous):

3x+9y=12 --> x + 3y = 4 after dividing by 3 5x-4y=3 ------------ x + 3y = 4 5x- 4y = 3 ----------- To use elimination, get the coefficients of either x or y to match Multiply the 1st eqn by 5: 5x +15y = 20 5x - 4y = 3 -------------- subtract 19y = 17 y = 17/19 ------ then sub for y in any eqn and find x.

OpenStudy (anonymous):

That help?

OpenStudy (anonymous):

Yep thanks